harrriet gave wilbur a 50 yd head start. how long did it take her to catch wilbur if her speed was 8 yds per second and his was only 6 yds per second?
Answer by stanbon(75887)

harriet gave wilbur a 50 yd head start. how long did it take her to catch wilbur if her speed was 8 yds per second and his was only 6 yds per second?
--------------
Wilbur DATA:
rate = 6 yds per sec. ; distance = x yds ; time = d/r = x/6 seconds
--------------
Harriet DATA:
rate = 8 yds per sec. ; distance = x+50 yds ; time = d/r = (x+50)/8 seconds
----------------
EQUATION:
time = time
x/6 = (x+50)/8
8x = 6x + 300
2x = 300
x = 150 yds.
------------
Harriet time = 150/8 = 75/4 = 18.75 seconds
===============
Cheers,
Stan H.